OpenStudy (anonymous):

How would you go about solving the following problem? lim x->0 [(x+1)^(1/3)-1]/x

OpenStudy (larseighner):

Okay, let's give this a shot. first, let's look at the numerator: (x+1)^(1/3) - 1 We are going to look at this as one of the factors of the difference of two cubes. Here is the formula for the difference of two cubes: a^3 - b^3 = (a - b)(a^2 + ab + b^2) You should stop now to see if this is sufficient information for you to solve the limit. =================================== that first factor a - b is what we are going to say is (x+1)^(1/3) - 1, or in other words, a = (x + 1)^(1/3) b = 1 Is that enough? So if we multiply the numerator and denominator by the second factor (a^2 + ab + b^2) that should not change anything. (we are multiplying by 1/1) ====================================== this looks nasty [(x+1)^(1/3) - 1] [(x+1)^(2/3) + (x+1)^(1/3)(1) + 1^2] / (x)[(x+1)^(2/3) + (x+1)^(1/3)(1) + 1^2] However the numerator is now [(x+1)^(1/3)]^3 - 1^3 That is where we got the formula from, remember. The numerator then is (x+1) - 1 = x x / (x)[(x+1)^(2/3) + (x+1)^(1/3)(1) + 1^2] The x's cancel, so we have 1/[(x+1)^(2/3) + (x+1)^(1/3)(1) + 1^2] The threat of division by zero being removed, we can solve by substitution Lim x->0 1/[(x+1)^(2/3) + (x+1)^(1/3)(1) + 1^2] = 1/[(0+1)^(2/3) + (0+1)^(1/3)(1) + 1^2] = 1/[1^(2/3) + 1^(1/3) + 1] = 1/3